How would you solve this: Laura and Dale drove in opposite directions. After two hours they were 68 kilometers apart. Laura drove 2 kilometers per hour faster than dale. How fast did Laura drive?
Answer by eschwartz(9)

Where D is the speed in km/h that Dale drives at,
2D+2(D+2)=68
Divide both sides by 2:
D+(D+2)=34km, or 2D+2=34
Subtract 2 from each side:
2D=32
Divide both sides by 2:
D=16
Now, since Dale drives at 16km/h, and Laura drives 2km/h faster, Laura drives at 18km/h.
In two hours, Dale dove 32km, and Laura drove 36km, for a combined total of 68km.